OpenStudy (anonymous):

I need help with the defferential equation: dy/dx = y/(y-x) , is it a linear D.E ?

OpenStudy (blurbendy):

http://tutorial.math.lamar.edu/Classes/DE/Definitions.aspx then look at definition 11

OpenStudy (anonymous):

tx but could you help me on this specefic problem?

OpenStudy (blurbendy):

can you write it in the form like definition 11 says?

OpenStudy (blurbendy):

I tried and I couldn't, so i don't think it's linear

OpenStudy (anonymous):

ok so what do you think this one is? Homogenoud? exact? seprable?

OpenStudy (anonymous):

It's not linear. Rewriting, you have \[(y-x)~\frac{dy}{dx}=x\] On the left, you have a function of the dependent variable multiplied by a power of its derivative. While I'm still here, it's not separable: there's no way to right \(\dfrac{x}{y-x}\) as a product of functions \(f(x)g(y)\). To determine whether it's homogeneous and/or exact, try writing it out in the form \[M(x,y)~dx+N(x,y)~dy=0\] If yes, it's homogeneous. Do you know how to determine whether an equation is exact?

OpenStudy (anonymous):

the dm/dy should be equale to dn/dx...it is not exact eighter

OpenStudy (loser66):

you can use substitute to solve it let y = Vx \[\frac{dy}{dx}=V +\frac{dV}{dx}V\] so, \[V+\frac{dV}{dx}=\frac{Vx}{Vx-x}=\frac{V}{V-1}\] \[\frac{dV}{dx}=\frac{V}{V-1}-V\] solve for V and then plug the solution back to substitute above to solve for y.

OpenStudy (anonymous):

\[\Large \frac{dy}{dx} =f(x,y)=f(tx,ty)\] Your differential equation fits the condition to be linear with the substitution method, a prior mentioned by @Loser66 already. Just consider to apply the initial test :-)

OpenStudy (loser66):

Thank you @Spacelimbus I take differential equation this summer, try to apply what I am taught. :)

OpenStudy (anonymous):

\[\Huge \checkmark \] Very good then (-: !